If a body covers distance x1 and x2 with velocity v1 and v2 respectively in Same time t .what is average velocity?

Grade:12th pass

1 Answers

talatamanjan
8 years ago
  1. given  a body covers x1 and x2 distances with velocities  vand vrespectively in equal time then the average velocity of that body is given by V=x1+x2/t1+t​ V=v1t+v2t/2t therefore V=v1+v2/2